OpenStudy (anonymous):

Excluding 28 itself, find the sum of the integral factors of 28

OpenStudy (prettygirl_shynice):

ok

OpenStudy (alexandervonhumboldt2):

do you know what is an interga factor?

OpenStudy (prettygirl_shynice):

yea

OpenStudy (alexandervonhumboldt2):

lets list all integer factors: 

OpenStudy (prettygirl_shynice):

The factors of 28 that are integer factors are: 1 and 28 2 and 14 4 and 7 If we're excluding 28 from the sum, we have: 1 + 2 + 4 + 7 + 14 = ?

OpenStudy (alexandervonhumboldt2):

factors are: 1, 2, 7, 2*7=14, 28, 2*2=4,

OpenStudy (anonymous):

1+2+4+7+14=28

OpenStudy (alexandervonhumboldt2):

no direct answers

OpenStudy (anonymous):

sorry did not know that rule

OpenStudy (anonymous):

:) remember that integral factors can be both positive and negative

OpenStudy (alexandervonhumboldt2):

yeah so also include same factors with a negative sign too

OpenStudy (anonymous):

ohhhhh